OpenStudy (anonymous):

Write the slope-intercept form of the line which passes through (4, 2) with a slope of 1/2.

OpenStudy (anonymous):

Equation of Line with given Slope and one point is: \[y - y_1 = m(x - x_1)\] Where, m= Slope = 1/2 And, \[(x_2, y_2) = (4,2)\] Put these values in the given equation and then solve to get the answer.. Able in doing or not??

OpenStudy (anonymous):

so would it be 1/2x+2?

OpenStudy (anonymous):

What you did in the first step?? Show me your work..

OpenStudy (anonymous):

Im still not that sure on how to do it

OpenStudy (anonymous):

See, first step to put the values in the given equaiton: We have, x1 = 4, y1 = 2, m = 1/2 Put these in equation: y - (2) = (1/2)*(x - 4) And now solve this equation to the simplest form: How can you do that?? By multiplying 2 both the sides, 2y - 4 = x - 4 Now solve and get the desired answer...

OpenStudy (anonymous):

Okay then would it be -2y=3 im not sure if i solved it quite right..... umm if i did then its still not right because the answers i can get are A = 1/2 x + 2 B. y = 1/2 x C. y = 1/2 x + 4 I thought it was a, but now im not sure

OpenStudy (anonymous):

See, slope intercept form is given by: y = mx + c So, try to reduce the equation I have solved above in this form: 2y = x So, y = (1/2)*x So, now what is the answer can you tell me??

OpenStudy (anonymous):

y=1/2x??(:

OpenStudy (anonymous):

Yes...

OpenStudy (anonymous):

Okay thank you so much!!

OpenStudy (anonymous):

Welcome dear..